OpenStudy (anonymous):

evaluate 2x^2 y^-1 when x=3and y=9

OpenStudy (anonymous):

You need to rewrite the expression so the exponents are positive \[\frac{ 2x ^{2} }{ y }\] then plug in 3 and 9 respectively \[\frac{ 2(3)^{2} }{ 9 }\] then you will do 3*3 and get 9*2 take that and divide by 9

OpenStudy (anonymous):

okay whats the answer and thank you

OpenStudy (anonymous):

just simplify \[\frac{ 18 }{ 9 }\]

OpenStudy (anonymous):

is it 2

OpenStudy (anonymous):

;) good job

OpenStudy (anonymous):

this is the compound interest questions are u good at it

OpenStudy (anonymous):

go for it

OpenStudy (anonymous):

you deposit $3,500 in an account that increases in value at a yearly rate of 5% find the balance after 12 years.

OpenStudy (anonymous):

give me a minute

OpenStudy (anonymous):

are u done

OpenStudy (anonymous):

Referencing this link: https://qrc.depaul.edu/StudyGuide2009/Notes/Savings%20Accounts/Compound%20Interest.htm plug into the formula to get \[3500(1+\frac{ 0.05 }{ 12 })^{12*12}\] This is assuming the interest is only applied once a year since it is not specified otherwise, you get \[3500(1+0.0041666666666)^{144}=3500(1.00416666666666)^{144}=3500*1.81984887\]

OpenStudy (anonymous):

what is the answer now

OpenStudy (anonymous):

just type 3500*1.8198 into a calculator

OpenStudy (anonymous):

ok answer this question four days before the SAT tests you studied 60 new vocabulary words. If you forget 5% of the words each day. How many words do you remember after 3 day

OpenStudy (anonymous):

Take 60*.95 that is after day 1 Take that answer *.95 that is after day 2 Take that answer *.95 that is after day 3 if you use a calculator just type the first thing, 60*.95 then hit *.95, *.95

OpenStudy (anonymous):

is the answer 51.4425